HUBBLE SPACE TELESCOPE DAILY REPORT #996 PERIOD COVERED: 0800 EST Daily Status Report as of 292/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 4892 (Duplicity Among Hyades Stars - Cycle 3 High) The FGSs were used to make astrometric observations of three possible binary stars VA722, VA292, and VA490 which is located in the Hyades open cluster. The OSS monitored the observation, and no anomalies were reported and observations looked normal. 1.2 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star systems HD41511 and HD60414 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these 1.2 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star systems HD41511 and HD60414 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. The observations of HD41511 occurred during LOS and could not be monitored in real-time. 1.3 Completed HSP/UV1/UV2/VIS/PMT/POL 5022 (HSP Photometric Calibration - Cycle 3) The High Speed Photometer was used to make calibration observations of star BD+75D325 to help calibrate the detectors/filter combinations being used in other subsequent proposals. Each observation was chosen to give about one million counts for each filter/detector combination. A total of ten observations were received. Only a few were monitored in real-time. The other observations remain on the science tape recorder and will be analyzed later. 1.4 Completed XSAFE (Safemode Recovery of the HST Science Instrument - HSP) The Safemode Recovery of the High Speed Photometer was executed as scheduled. No problems were reported. 1.5 Completed FOC 5043 (FOC Internal Relative Calibrations) The Faint Object Camera (f/96) was used to measure the pixel to pixel response by using the internal LEDs. The OSS monitored the observations, and the data quality was assessed. No anomalies were reported. 1.6 Completed FOS 5049 (FOS Cycle 3 Wavelength Calibration: Internal/External Offsets) The Faint Object Spectrograph was used to make observations of external object NGC-6833 and also of the internal lamp to determine the offset for pixel to wavelength calibration. Object NGC-6833 was used because it lacks sufficient emission lines for grating G130H. The observations were executed as scheduled, and no problems were reported. 1.7 Uplinked Updated Ephemeris Tables, Per MOSES 2871 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Significant HSTARs: None 2.4 Operations Note: Internal Simulation Type 10 (EVA#4) is being conducted today, Tuesday Oct. 19,1993 from 8AM until 8PM. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of GHRS/2 5251 (WF/PC Shutter Control AP Test) 3.2 Complete GHRS/2 5041 (GHRS Sensitivity Monitor - Cycle 3) 3.3 Complete Two Sets of FOC/96 4445 (SINS: The Supernova Intensive Study - 87A Revisit - Cycle 3 High) 3.4 Complete WFC 4497 (Parallel High Resolution Imaging of Defuse Objects in Magellanic Clouds) 3.5 Complete FGS 4938 (Parallaxes of Hyades cluster Members - Cycle 3 High) 3.6 Complete WFC 4848 (GTO Low Lattitude Parallel Program _ Cycle 3 High) 3.7 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.8 Complete PC 4308 (Snapshot Survey for Companions to Planetary Nebula nuclei - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-